3 years ago I was half the age. I will be 7 years from now how old am I
Answer by Cromlix(4381)

x - 3 = 1/2(x + 7)
x - 3 = 1/2x + 7/2
Collect like terms:
x - 1/2x = 7/2 + 3
1/2x = 13/2
Cross multiply
x = 13
I am 13 years old.
Hope this helps.
:-)